OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge \frac{d}{dx} \sin^2 (e^{\sin^2 t})\]

OpenStudy (anonymous):

is that \[dy/dt\] or with respect to x

OpenStudy (lgbasallote):

so im thinking \[\large 2\sin (e^{\sin ^2 t}) \cdot \cos (e^{\sin ^2 t} \cdot e^{\sin^2 t} \cdot 2\sin t \cos t\]

OpenStudy (lgbasallote):

is that right?

OpenStudy (lgbasallote):

\[\large 4\sin (e^{\sin^2 t}) \cos (e^{\sin ^2 t}) e^{\sin ^2 t} \sin t \cos t\]

OpenStudy (lgbasallote):

@Jonask dy/dx

OpenStudy (anonymous):

right

OpenStudy (lgbasallote):

ahh glad to hear that